VSEPR, Hybridizations, aproximate values, bond angle

What hybridization would you expect for the {rm C} and {rm N} atoms in urea according to the VSEPR model.

McMurry.Ch07.035.jpgWhat approximate values would you expect for the various bond angles?N-C-O angle=N-C-N angle=C-N-H angle=H-N-H angleWhat are the actual hybridizations based on the molecular model shown?
{rm C} has an sp^3hybridization, {rm N} has ansp^2 hybridization.
{rm C} has an sphybridization, {rm N} has ansp^3 hybridization.
{rm C} has an sp^2hybridization, {rm N} has ansp^3 hybridization.
{rm C} has an sp^2hybridization, {rm N} has ansp^2 hybridization
What are the bond angles based on the molecular model shown?N-C-O angle=N-C-N angle=C-N-H angle=H-N-H angle=

Could you break it down steps for me to understand how arrive to the conclusion, including clear formulas.Not just summarizing it. Would be helpful if you break it down in detain as I would like to understand it. Thanks. Make it a clearformula
Answer #1


Here in the structure C is in sp2 hybridization and the two N's are in sp3 hybridization.

N-C-O bond angle = 1200N-C-N bond angle = 1200C-N-H bond angle = 1070H-N-H bond angle = 1070
answered by: Jazmine Beatty